Earl Thomas has received a raise from $50,000 a yr to $55000 a yr . What was Earls raise stated as a percentage? Inflation averaged 3.3% for the yr. What was earls real income after the raise? What was Earls real raise stated as a percentage ?

For the first question my answer was 10
My second answer is 53242
I am stuck in the 3rd question. I tried 55000-53242=1758 then 1758%55000x100=3.19

Answer by math_tutor2020(3821)


Problem 1

difference = new salary - old salary
difference = 55,000 - 50,000
difference = 5,000

percent change = (difference)/(original)
percent change = (5,000)/(50,000)
percent change = 1/10
percent change = 0.10
percent change = 10%

Your answer for problem 1 is correct.
Earl got a 10% raise if we ignore the aspect about inflation.

---------------------------------------------------

Problem 2

Let's say Earl wants to buy a car worth $50,000
Without inflation, the price would stay where it is.

With the 3.3% inflation, the price increases to 1.033*50,000 = 51,650

Effectively Earl loses $1,650 in earnings due to this rate of inflation (since 51,560-50,000 = 1,650)
In other words, he would need to earn an additional $1,650 in order to keep up with inflation.
It is said that the inflation "eats" away at those earnings so to speak.

His new salary of $55,000 is really 55,000 - 1,650 = 53,350 dollars when considering inflation.

Notice that without inflation, we could say
50,000*1.10 = 55,000
and with the 3.3% inflation, we could say
50,000*(1.10-0.033) = 53,350

The 10% raise is really 10% - 3.3% = 6.7% when considering inflation.

Answer: $53,350

---------------------------------------------------

Problem 3

Earlier we calculated the real raise percentage was 6.7% (refer to the tail end of problem 2).

Here's another approach

difference = new salary - old salary
difference = 53,350 - 50,000
difference = 3,350

percent change = (difference)/(original)
percent change = (3,350)/(50,000)
percent change = 0.067
percent change = 6.7%

Answer: 6.7%